A window consists of a rectangular piece of clear glass with a semicircular piece of
colored glass on top; the colored glass transmits only 1/2 as much light per unit area as the the clear
glass. If the distance from top to bottom (across both the rectangle and the semicircle) is 2 meters and
the window may be no more than 1.5 meters wide, find the dimensions of the rectangular portion of the
window that lets through the most light.
1
Expert's answer
2020-07-29T14:39:31-0400

Let 𝑟 be a radius of a semicircle, "H" be a distance from top to bottom. Then the horizontal side of the rectangle is equal 2𝑟, vertical side of the rectangle is equal 𝐻 − 𝑟. Effective area of the window is


"S=S_{rect}+kS_{semi}=2r(H-r)+k\cdot{1\over 2}\pi r^2"

Given


"H=2\ m, k={1\over 2}, 0<2r\leq 1.5\ m"

Substitute


"S=S(r)=2r(2-r)+{1\over 2}\cdot{1\over 2}\pi r^2"

"S=S(r)=4r-2r^2+{1\over 4}\pi r^2, 0<r\leq0.75"

Find the critical number(s)


"S'=(4r-2r^2+{1\over 4}\pi r^2)'=4-4r+ {1\over2}\pi r"

"S'=0=>4-4r+ {1\over2}\pi r=0""r={8\over 8-\pi}"

First derivative test

If "0<r<\dfrac{8}{8-\pi}, S'>0, S" increases.

If "r>\dfrac{8}{8-\pi}, S'<0, S" decreases.

The function "S(r)" has a local maximum at "r=\dfrac{8}{8-\pi}."

Since the function "S" has the only extremum, then the function "S" has the absolute maximum "r=\dfrac{8}{8-\pi}."


"\dfrac{8}{8-\pi}>1.5"

Hence we have to take "r=0.75 m"

The width of the rectangular portion of the window  is "2r=1.5m."

The height of the rectangular portion of the window  is "2-0.75m=1.25m"

The dimensions of the rectangular portion of the window that lets through the most light are


"1.5m\times1.25m"